40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 138 of 330  Not logged in ELOG logo
ID Date Author Type Category Subject
  9684   Mon Mar 3 11:55:39 2014 KojiUpdateCDSfb timing was off

We need to correctly setup crontab or rc.local for the frontend machines.

  9683   Mon Mar 3 10:42:53 2014 JenneUpdateCDSfb timing was off

...yet again.

lsc and sus needed mxstream restarts after I restarted the ntp on fb.

  9682   Thu Feb 27 22:25:29 2014 ranaUpdateSUSOplev Tuning Party - round 1 commentary

  in order to Win in Loop Tuning, you must draw a cartoon of the cost function on the whiteboard before starting. Some qualitative considerations from our Workshop:

  1. We want to use the oplev servo to reduce the motion of the mirror in the frequency band where the Oplev is quieter than the mirror, w.r.t. inertial space.
  2. We can estimate the true mirror motion by some simple stack / pendulum model and compare it to the Oplev noise (not the dark noise). There are several contributions to the mirror angular motion due to the cross-coupling in the stacks and pendula.
  3. Below ~0.2 Hz, we think that the oplev is not the right reference, but this is not quantitative yet.
  4. The high frequency noise in the OPLEV ERROR is definitely electronics + shot noise.
  5. We cannot increase the gain of the loop without posting some loop measurements (Bode + steps). Also have to post estimates of how much PRCL noise is being introduced by the Oplev feedback. Oplev feedback should make less length noise than what we have from seismic.

Give us a cost function in the elog and then keep tuning.

  9681   Thu Feb 27 13:11:13 2014 steveUpdatesafetysafety audit correction

Quote:

 

 We had our annual safety inspection today.  Our SOPs are outdated. The full list of needed correction will be posted tomorrow.

 

The most useful found was that the ITMX-ISCT ac power is coming  from 1Y1 rack. This should actually go to 1Y2 LSC rack ?

 Please test this so we do not create more ground loops.

 Linus-1, Nodus and others  ac cords can be moved over to new blank yellow extension cord with multiple recepticals.

 Remove two red extension cords going to Smart UPS

Attachment 1: noDaisyChaining.jpg
noDaisyChaining.jpg
Attachment 2: ZbThumbnail.info
  9680   Thu Feb 27 01:02:57 2014 JenneUpdateSUSOplev Tuning Party - round 1

[Jenne, Vivien]

We had an oplev tuning party this afternoon.  What we have learned is that we don't have a lot of intuition yet on tuning loops.  But, that was part of the point - to build some intuition. 

I took responsibility for the PRM, and Vivien took ITMX.  I think, in the end, all changes were reverted on ITMX, however Vivien took some data to try and make a computer-generated controller.  Before we got started, I locked and aligned the PRMI, and we centered the PRMI-relevant oplevs.

I moved my "boost bump" around a bit, to do more at higher frequencies, but had to sacrifice some of the "oomph", since it was starting to eat up too much phase at my UGF of ~8Hz.  I also made the stack resonant gain higher Q and lower height so that it didn't eat so much phase.  In the end, I have 25 degrees of phase margin, which isn't really great, but I do win a factor of 2 around 2 and 3 Hz.  Also, now I'm able to engage the 3.2 resgain at all, whereas with the previous filter shape I was not able to turn it on.

PRM_oplevTuning_26Feb2014.pdf

Maybe it's because I really want it to have helped, but I feel like the POP spot isn't moving as much when I'm locked on PRMI sidebands as it was earlier (we were seeing a lot of low frequency (few Hz) motion).  So, I think I did something good.

  9679   Wed Feb 26 23:14:07 2014 JenneUpdateCDSfb timing was off

....fb timing issue happened again.

I thought that it was the thing that Koji and I saw the other day, where it was individual front end computers that had lost ntp sync, since it wasn't every core on every computer that was red, but reconnecting to the ntp server on c1lsc didn't do anything.  I then tried reconnecting to the ntp server on fb, and that fixed things right up.  Annoying.

  9678   Wed Feb 26 10:08:14 2014 SteveUpdateIOOIOO trend

 

 The MC is happy (but only for this tiny snapshot in time and most probably will go dysfunctional again as it has been for several months, as of this writing)

Attachment 1: IOOtrend3&24h.png
IOOtrend3&24h.png
  9677   Wed Feb 26 02:20:35 2014 JenneUpdateIOOMC unhappy

I've asked Manasa and Q to have a look at the MC in the morning.  Rana and I have found it to be slightly uncooperative in relocking after a lockloss.

The concern is that we may be (by actuating on things during lock, or during a lockloss) ringing up some mode, maybe a violin mode in one of the suspensions, maybe a PZT mode of some sort.  If we are, and then we have to push with the PZT on the laser to lock things, that may be why the laser's PZT RMS (on the FSS screen) is so often above 1Vrms.  When we close the PSL shutter, the rms is low, like 0.6 or something, and it stays flat.  As we've all see many a' time, the red trace on the top projector plot is pretty erratic throughout the day when the MC is locked or trying to lock.

We have found that just letting the autolocker go doesn't seem to work very well, and sometimes the MC just doesn't want to re-lock.  Closing the PSL shutter or disabling the autolocker for a few minutes (5ish) doesn't do anything, but leaving it closed for a long time (30 ish minutes) helps a lot.  The MC  will relock immediately after a nice long break. 

 

  9676   Wed Feb 26 01:49:08 2014 JenneUpdateLSCChanging PRCL offset changes REFL 165 degeneracy

I have measured the sensing matrix at a variety of PRCL offset values.

DemodPhaseSeparation.pdf

During this each measurement, I also took a 20 second average of the POP 2f signals and the ASDC signal:

POP_AS_PDvalues.pdf

All of this data was taken during a single lock stretch. 

If / when I do this again, I want to go out to larger offsets.  I won't take as many points, but I do want to see how far I can go before I lose lock, and what the phase separation looks like at larger offset values (this time, I stopped at +700 counts which is about 0.7nm, to start checking the negative values. MC has been unhappy, so I wasn't able to take very many negative offset values.) 

I conclude that these sensing matrix measurements do see changes in the phase separation with PRCL length offset (what we saw / said yesterday), but that they do not line up with Q's simulation from this afternoon in elog 9671.

The simulation says that we shouldn't be seeing large phase changes until we get out to several nanometers, however the measurement is showing that we get large phase chnages with picometer scale offsets.  Yesterday, Rana and I said that the offsets due to RAM were small (of order picometer), and that they were therefore likely not important (elog 9668).  However, now it seems that the RAM is causing significant length offsets which then cause poor MICH/PRCL phase separation.

To Do List:

* Confirm MIST simulation with Optickle.

* Look at sensing matrix data pre-lockins (in the raw sensors).

* Check that there is no clipping anywhere in the REFL path (at least out of vacuum), and that the beam is sufficiently small on all 4 REFL diodes.

* Calculate the new PRC g-factor with the new length.

  9675   Tue Feb 25 23:38:05 2014 rana, jenneUpdatePEMGUR1 Z channel excess noise: oscillating Z channel

Last night we noticed an excess in the GUR1Z seis BLRMS on the StripTool. It was in the 0.1 - 0.3 Hz band. The rumor in the control room was that "this kind of noise has been showing up at night recently".

AS it turns out, this was not some environmental noise around the 40m at night, but instead its some internal servo oscillation in the GUR1 Z channel. In the Guralp seismometers, each channel is a different mechanical sensor (unlike the STS or T240), so when a single channel gets noisy it doesn't always implicate the others.

My guess is that the oscillation came from the Z channel needing to be recentered. I power cycled the interface box just now. The oscillation had already gone away, but I thought this might reduce the excess noise. Maybe it did, but the effect is tiny. You can see in the oscillation reference that the low frequency noise is high, but in the new trace its still kind of high. Needs to be re-centered correctly with the paddle. Or add a centering button to the interface box.

Attachment 1: gur1z.pdf
gur1z.pdf
  9674   Tue Feb 25 18:16:22 2014 JenneSummaryLSCEven more violin filters

A new violin mode at 1303 Hz was ringing up this afternoon.  Rana and I added a notch for this.

RXA: while the mode at 1303.6 Hz was ringing down, I used the narrowband DTT technique to measure the Q (after turning on the notch in SUS-PRM_LSC). So its another frequency in the PRM (not the BS).

The time that it takes for 2 -foldings is 652 s, which implies that Q = pi*f*tau = 1.3e6. This seems too high by a factor of ~10, so my guess is that there is still some feedback path happening. The previous bandstop filter was centered around 1285 Hz and seems also weird that the PRM would have 2 violin modes with such different frequencies. Is the mirror rotated around the optic axis such that the standoffs are not at the same height?

Attachment 1: PRMvio2.png
PRMvio2.png
  9673   Tue Feb 25 17:27:41 2014 JenneUpdateLSCREFL signals calibrated

I have recalibrated the REFL signals.

I first adjusted the demod phases until the I-signals lined up with the I-phase in the sensing matrix plot:

SensMat_25Feb2014.png

I then balanced the ITM drives by pushing on -1*ITMX and +1.015*ITMY, and seeing a minimum of MICH actuation in the I-phase of REFL55 (the PD I was locking with).

I then took a nice long measurement with DTT, and measured the peak heights in I and Q for each REFL diode.  I was driving PRM with 100 cts at 675.1Hz, and ITMX with 1000 cts at 452.1 Hz (and matching ITMY drive, to make pure MICH).  Knowing these numbers, and the actuator calibrations (PRM elog 8255, ITMs elog 8242), I know that I was driving PRCL by ~4.3 pm, and MICH by ~23 pm. 

For the I-phase calibrations, I find the peak height at the PRCL drive frequency, and divide 4.3 pm by that height.  For the Q-phase calibrations, I find the peak height at the MICH drive frequency, and divide 23 pm by that height.

This gives me the following calibrations:

  Calibration [picometers / count]
REFL 11 I    0.15
REFL 11 Q   21.6
REFL 33 I    1.06
REFL 33 Q  209
REFL 55 I    0.9
REFL 55 Q   27       
REFL 165 I    0.1
REFL 165 Q   11.6

 My calibrated REFL spectra then looks like:

Calibrated_25Feb2014.pdf

  9672   Tue Feb 25 16:54:57 2014 steveUpdatesafetysafety audit 2014

 

 We had our annual safety inspection today.  Our SOPs are outdated. The full list of needed correction will be posted tomorrow.

 

The most useful found was that the ITMX-ISCT ac power is coming  from 1Y1 rack. This should actually go to 1Y2 LSC rack ?

 Please test this so we do not create more ground loops.

  9671   Tue Feb 25 16:07:33 2014 ericqUpdateLSCChanging PRCL offset changes REFL 165 degeneracy

 And glossing over the MICH offset, here's the PRC offset plots in displacement, rather than radians.

The simulation is actually slightly different now. I now use nominal ITM T values (T=.014) instead of the random R=.99 I had in place. 

MICHvPRCLangle_wOffset.pdfMICHvPRCLangle_wOffset_fullscale.pdf

(correction: Field Power should be Field Amplitude in the first plot)

  9670   Tue Feb 25 14:48:49 2014 ericqUpdateLSCChanging PRCL offset changes REFL 165 degeneracy

After speaking with Jenne and Gabriele, I did a little bit of simulating based on my earlier code that looked at the angle of MICH vs. PRCL, just with cavity detuning instead of macroscopic length change.

The zero point in the following plots is with the PRC locked on the sideband. The PRC detuning was done by changing the PRM-BS microscopic length (in terms of phase), and the MICH detuning was done by adding half of the detuning to the BS-ITMY distance, and subtracting half of it from the BS-ITMX distance. 

MICHvPRCLangle_wOffset.pdf

 

This plot is in terms of radians, so to roughly relate it to line width, here's a plot of the POP powers as a function of the PRC detuning. 

SBprclPeaks.pdf 

  9669   Tue Feb 25 02:46:38 2014 rana, jenneUpdateLSCChanging PRCL offset changes REFL 165 degeneracy

[Jenne, Rana]

We put offsets in the PRCL and MICH loops, and measured sensing matrices for each condition. 

What we found was that PRCL offsets of order 1/20th a linewidth (calibration to be checked tomorrow) would give significant changes in the angles of the REFL signal sensing matrix elements.  We broke MICH lock before we were able to put in a significant enough offset to see the demod phases change.

Because there are so many plots, I've put them together in a pdf. Each page has a set of radar plots for sensing matrix elements.  On the bottom of each page I note what our MICH and PRCL offset values were, and where the data is saved (in the 40m scripts directory). To see the differences, make sure your pdf viewer is set to single-page, not scrolling.

PRC_offsetCheck_24Feb2014.pdf

One major thing that we noted was that putting in a PRCL offset also changed the MICH offset.  When we increased the PRCL offset, we saw the AS port get brighter (but not as bright as when we were putting in large MICH offsets). 

Tomorrow, I need to check the calibrations we were using, to see how many meters we were moving the optics.  Also, Q, Gabriele and I need to meditate and do some modelling to figure out why the length offset could be affecting the degeneracy so strongly. 

  9668   Tue Feb 25 00:00:01 2014 rana, jenneUpdateLSCreasons that the REFL signals may be degenerate now

We're exploring some effects which may give some funny macroscopic detuning and cause a near phase degeneracy in the REFL RF signals (see radar plot from Jenne below).

1) Alignment: we centered the oplevs to reduce fluctuations and then tweaked the BS and PRM alignment to build up the power. No significant change in the RF phases of the DOFs.

2) Measuring RAM: we set the dark offsets (by hand since the Masayuki script doesn't really work well anymore) to with 1 counts. We then locked the MC, misaligned the ITMs, and looked at the REFLOUT16 channels using the following command line:

z avg 12 C1:LSC-REFL11_I_OUT16 C1:LSC-REFL11_Q_OUT16 C1:LSC-REFL33_I_OUT16 C1:LSC-REFL33_Q_OUT16 C1:LSC-REFL55_I_OUT16 C1:LSC-REFL55_Q_OUT16 C1:LSC-REFL165_I_OUT16 C1:LSC-REFL165_Q_OUT16
C1:LSC-REFL11_I_OUT16     -12.04
C1:LSC-REFL11_Q_OUT16     -14.34
C1:LSC-REFL33_I_OUT16       0.43
C1:LSC-REFL33_Q_OUT16      -0.28
C1:LSC-REFL55_I_OUT16       2.84
C1:LSC-REFL55_Q_OUT16       5.64
C1:LSC-REFL165_I_OUT16      4.40
C1:LSC-REFL165_Q_OUT16      0.10

So these offsets are small in counts. In meters this corresponds to....less than 3 pm for any of the I signals.

Refl11I = 2.06e-12 meters

Refl11Q = 2.94e-10 meters

Refl33I = 5.28e-13 meters

Refl33Q = 1.07e-11 meters

Refl55I = 2.71e-12 meters

Refl55Q = 3.55e-11 meters

Refl165I = 3.07e-13 meters

Refl165Q = 8.63e-14 meters

 

 

3) Next we want to put large offsets into the error points of the loops

4) Change modulation depth

5) Check IMC length (todo for Q/Manasa for Tuesday - Wednesday)

  9667   Mon Feb 24 23:43:10 2014 ranaSummaryGeneralToDo

1) Fixup REFL165: remove ND filters, get box for PD, dump diode reflections, put less light on diode, change DC transimpedance (?), max power dissipation on BBPD < 0.5 W w/ 25 V bias. Perhaps replace OP27 with TLE2027.

2) Make plan for fixing fiber layout up and down the arms. Need tubing for the whole run. Don't make it cheesy. Two fibers per arm.

3) Fix LSC model to allow user switching of whitening. Get back to working on AutoLock scripts (not Guardian).

4) Manasa, Q, Jenne, tune Oplev servos Tuesday morning/afternoon.

5) Reconnect the other seismometers (Steve, Jenne). For real.

6) Balance PRMI coils at high frequency.

  9666   Mon Feb 24 17:59:31 2014 RANAUpdateElectronicsMeasured REFL165 demod board

 

 Demod boards should be at 90 deg, not 82.7 or 12 or yellow or ****. We should re-inject the RF and then set the D Phase in the filter module to make the signals orthogonal. 165 is a challenging one to get right, but its worth it since the signals are close to degenerate already.

  9665   Mon Feb 24 17:21:42 2014 SteveUpdateGreen Lockinggreen fiber status today

Quote:

Alex, Gautam and Steve,

Single mode fiber 50m long is layed out into cable tray that is attached to the beam tube of the Y arm.

It goes from ETMY to PSL enclosure. It is protected at both ends with " clear- pvc, slit corrugated loom tubing " 1.5" ID

The fiber is not protected between 1Y1 and 1Y4

 The X -arm fiber is in the high cable tray and it has has  coupler mounts.

 The Y -arm fiber is in the low cable tray and it has no coupler mounts.

 The fibers are only protected at entering and exiting the trays.

 We have only 68 ft spare 1.5"  ID protective plastic tubing.

Attachment 1: etmy_F@1Y2.JPG
etmy_F@1Y2.JPG
Attachment 2: etmy-F@PSL_.jpg
etmy-F@PSL_.jpg
Attachment 3: etmx_F@se.JPG
etmx_F@se.JPG
Attachment 4: etmx_F@1Y8.JPG
etmx_F@1Y8.JPG
Attachment 5: etmx_F@PSL.JPG
etmx_F@PSL.JPG
Attachment 6: etmy_F@ee__.jpg
etmy_F@ee__.jpg
  9664   Mon Feb 24 16:26:14 2014 JenneUpdateCDSNTP fell out of sync on front end machines - fixed

[Koji, Jenne]

Koji noticed that the time on the front-end detail screens was not correct, and that the GPS time was not matching up between different models.  Koji ran the following on all front-end machines, and on nodus:

sudo ntpdate -b -s -u pool.ntp.org

Now, everything is fine, and every status light on the cds overview screen is green.

  9663   Mon Feb 24 15:25:29 2014 JenneUpdateCDSComputer weirdness with c1lsc machine

The LSC machine isn't any better, and now c1sus is showing the same symptoms.  Lame.

The link lights on the c1lsc I/O chassis and on the fiber timing system are the same as all other systems.  On the timing card in the chassis, the light above the fibers was solid-on, and the light below blinks at 1pps. 

Koji and I power-cycled both the lsc I/O chassis, and the computer, including removing the power cables (after softly shutting down) so there was seriously no power.  Upon plugging back in and turning everything on, no change to the timing status.  It was after this reboot that the c1sus machine also started exhibiting symptoms. 

  9662   Mon Feb 24 13:40:13 2014 JenneUpdateCDSComputer weirdness with c1lsc machine

I noticed that the fb lights on all of the models on the c1lsc machine are red, and that even though the MC was locked, there was no light flashing in the IFO. Also, all of the EPICS values on the LSC screen were frozen.

Screenshot-Untitled_Window-1.png

I tried restarting the ntp server on the frame builder, as in elog 9567, but that didn't fix things.  (I realized later that the symptom there was a red light on every machine, while I'm just seeing problems with c1lsc. 

I did an mxstream restart, as a harmless thing that had some small hope of helping (it didn't). 

I logged on to c1lsc, and restarted all of the models (rtcds restart all), which stops all of the models (IOP last), and then restarts them (IOP first).  This did not change the status of the lights on the status screen, but it did change the positioning of some optics (I suspect the tip tilts) significantly, and I was again seeing flashes in the arms.  The LSC master enable switch was off, so I don't think that it was trying to send any signals out to the suspensions.  The ASS model, which sends signals out to the input pointing tip tilts runs on c1lsc, and it was about when the ass model was restarted that the beam came back.  Also, there are no jumps in any of the SOS OSEM sensors in the last few hours, except me misaligning and restoring the optics.  I we don't have sensors on the tip tilts, so I can't show a jump in their positioning, but I suspect them.

I called Jamie, and he suggested restarting the machine, which I did.  (Once again, the beam went somewhere, and I saw it scattering big-time off of something in the BS chamber, as viewed on the PRM-face camera).  This made the oaf and cal models run (I think they were running before I did the restart all, but they didn't come back after that.  Now, they're running again).  Anyhow, that did not fix the problem.  For kicks, I re-ran mxstream restart, and diag reset, to no avail.  I also tried running the sudo /etc/init.d/ntp-client restart command on just the lsc machine, but it doesn't know the command 'ntp-client'. 

Jamie suggested looking at the timing card in the chassis, to ensure all of the link lights are on, etc.  I will do this next.

  9661   Mon Feb 24 13:21:00 2014 JenneUpdateElectronicsMeasured REFL165 demod board

I measured the REFL 165 demod board's I/Q separation. 

Our 11MHz signal is currently 11.066092 MHz, so I put a signal to the RF input of the REFL165 demod board at 165.992380 MHz (15*11 MHz + 1kHz), with a signal of -13 dBm.

I then used the SR785 to measure the transfer function between the I and Q output channels. 

I got 82.7 degrees, at -0.64 dB. (I don't remember now if I had I/Q, or Q/I, not that it really matters). So, it seems that the REFL165 demod board has good separation, and at least isn't totally broken.

  9660   Fri Feb 21 12:45:57 2014 ericqUpdateLSCEquivalent Displacement Noise from QPD Dark Noise in SQRTINV

EQ UPDATE: Measured it wrong the first time, fixed now.

I measured the spectra of the SQRTINV channels from dark QPDs, with offsets adjusted to imitate various transmission levels. (While the dark noise stays constant in terms of, say, TRX counts, 1/sqrt(TRX) isn't linear, and so the noise coupling depends on the TRX offset). 

SQRTINVspectra.pdf

I did some calculations to turn this into the equivalent displacement noise when using SQRTINV as an error signal. This depends on where on the fringe you are locking, since the slope of SQRTINV vs. position is not constant, and can only really be treated as linear down to about 1/3 of a line width away from full resonance. In my calculations, I assumed a coupled arm line width of 38pm, and a full transmission of 700 counts in TRX/Y. 

The QPD dark noise RMS when two line widths away (TR = 40) is about 5fm, and only goes down from there. 

SQRTINV_DarkNoise.pdf

  9659   Wed Feb 19 22:47:26 2014 JenneUpdateLSCALS locked using LSC model, Common & Diff transitioned to IR transmission signals

[Jenne, Koji, Manasa, EricQ]

Today we successfully locked the ALS using the LSC system, with filters that are good for both the IR PDH and the ALS locking.  We tried PRFPMI, but were unable to hold PRMI lock while the arms were held with ALS.  We combined the ALS signals into common and differential signals, and successfully transitioned over to a combined set of 1/sqrt(TRANS) signals for the common mode part of the lock (differential stayed with ALS). 


Locking the ALS using filters in the LSC system that are also good for IR PDH

The biggest difference between the ALS and LSC filters were the ones used for lock aquisition. At Koji's suggestion, I made FM5 of the LSC servos (for X and Y arms) the filter needed for ALS locking.  Then, I made FM4 into a combination of old LSC FM4 and FM5, as well as an inverse of the new FM5, so that when both FM4 and FM5 are engaged, the servo shape is the same as the old LSC.  I left the other LSC filters where they were.  I replaced the FM1 +6dB with the combined integrators (really, just gentle DC boosts) for the ALS, since we were never using this +6dB filter module.  The LSC resonant gain filter for the bounce mode also included a resgain for 18.5 Hz.  I don't know what that was for, and it was eating into phase that I needed, so I removed it.

The other filter that changed significantly was the Boost filter.  The ALS system had been using more DC gain than the LSC had.  However, the current ALS boost filter (in FM10 of the old ALS servos) was eating too much phase near my UGF.  So, I scooted the whole boost filter to lower frequencies, to give myself some extra phase margin.  The boost was set to "zero history", "zero crossing", with 0.01 tolerance and an 8 second timeout.  Setting it to zero crossing with a low tolerance, rather than just ramping it on, was the key to engaging the boost.

ALS_newVSoldBoosts.pdf

I had to be so careful about phase margin, since I lost ~15 degrees of phase at 200 Hz from the lag of going through the RFM network.  This was pretty frustrating, but I don't have a better plan yet, save moving the c1als model and ADC to the SUS machine, which has Dolphin access to the LSC.  I may back off my safety margin, and give myself some gain in the boost back at 10Hz, since we are now seeing too much noise at 10Hz in the closed-loop spectra.  I also "cheated" and lowered my UGF from the ~150Hz it used to be in the ALS model, to 100Hz, where I was closer to the top of the new phase bubble.

With the new filter situation, I was able to lock the Xarm (the one I was using for design work) with both IR and ALS.  To lock IR, the "restore" script still works. For the ALS, we should put in a separate "restore" script into the IFO_CONFIGURE screen. 

The ALS locking procedure is as follows:

* Prepare ALS and green locking.  Green locked to 00 mode, alignment all nice, etc, etc.  Beatnote within 100MHz on spectrum analyzer.  If doing both arms, try to get beatnotes on opposite sides of PSL, to keep crossbeatnotes at higher frequencies, and out of the way.

* Turn on Watch script.

* Set LSC parameters (this is where a new restore script will come in handy): 

       * Zeros in RFPD columns of input matrix (i.e. POX and POY).

       * Ones in AUX input matrix elements.

       * Zeros in power normalization matrix rows for arms.

       * All FM triggers for arms set to "Man" for manual.

       * Override main trigger, so that signals are always going through to the servo.

       * Only FM5 engaged in arm servo.

       * Gain of servo set to zero, output on, then engage main LSC master switch.  ETM output on.

* Clear history in phase tracker.

* Check sign of gain using + or - 0.1 in the servo.  You'll know if you got it wrong (the ETM will be kicked, and the beatnote will fly around).  If you didn't get it wrong, you probably got it right.

* Increase gain to about 12 (with correct sign).

* Engage FM1 (gentle DC boost), FM6,7,8 (resonant gains for stack, bounce, roll)

* Wait a few seconds for filters to settle, then engage FM9 (boost).

* Run find IR resonance script.

* Move off resonance by ~36 counts (12 times the +3 script).  This number comes from trying to be completely off the IR resonance, even when the PRMI was locked.

* Do whatever locking (ex. PRMI) you set out to do.


 PRFPMI attempt

After locking both arms with ALS using the LSC system, we attempted to lock the PRMI.  We were able to lock PRMI on REFL55 I&Q, REFL33 I&Q, and REFL55 I&AS55Q before the arms were locked, so we were hoping that we wouldn't have too much trouble.

We found the IR resonance for both arms, then moved off resonance.  Then, restored the PRM.  For REFL55, Koji coarsely turned the REFL 55 demod phase from 16 degrees to 87, while we were locked on the carrier.  After this, I stepped farther and farther from the IR resonance, since at first I found that our transmitted powers were something like 4, rather than almost zero, so the demod phase may not be totally correct.  

We were having trouble, so we locked the PRMI on carrier using REFL55 I and AS55 Q, with 1's in both elements in the input matrix.  MICH gain was about -10, PRCL +0.010.  We used this time to tweak up the alignment of the PRMI.  At some point, Koji tweaked the REFL33 demod phase from 124 to 134 degrees.  Then we switched back to sideband locking.  After some trials with REFL55 I&Q, and REFL55/AS55, we went to REFL33 I&Q.  REFL33I->PRCL was 1.556 in the input matrix, and REFL33Q->MICH was -0.487.  No other elements in the input matrix.  MICH gain was reduced to -6, PRCL gain to -0.020.  MICH FMs 3,6,9 triggered, PRCL FMs 2,3,6,8,9 triggered.  We were able to keep short locks on the order of ~10 seconds, but not longer. We played with every parameter we could think of (alignment being good is one of the most important!), but were not able to keep better lock.  The POP spot is moving around a lot, so the PRCL ASC needs to be examined, hopefully tomorrow.

We started losing the Xarm lock fairly regularly, I'm not sure why, but the Yarm was locked for almost 2 hours straight, held off resonance with ALS!


 ALS Common and Differential, transition to IR control

We set PRMI aside for the rest of the night, and looked at using ALS to control the arms in common and differential modes. 

Regular ALS locking procedures were used (see above), with the exception of the AUX input matrix:

  1/sqrt(TRX) 1/sqrt(TRY) ALSX ALSY
XARM (common) 0 0 +1 -1
YARM (differential) 0 0 +1 +1

 Since the beatnotes were on opposite sides of the PSL frequency, the common and differential modes look opposite of what you'd expect. 

We then used the regular find IR resonance scripts running simultaneously, which worked really well to find both arms' IR resonance points.

I put a 1 count offset in the Xarm servo (which was our proxy for common mode), although in retrospect this should have been +0.5 in ALSX, and -0.5 in ALSY, so that our signals going through the input matrix were at their zero crossings.  Anyhow, this offset put us at about half fringe on both arms (transmissions were about 0.6). 

Koji set the offsets in the 1/sqrt(trans) filter banks before the input matrix so that they would have zero crossings at this point (avg the IN1, put negative of that value into the offset). 

We then stepped the input matrix values until our common mode (Xarm) row was:

  1/sqrt(TRX) 1/sqrt(TRY) ALSX ALSY
XARM (common) -0.7 -0.7 0 0

We left the differential (YARM) row alone, so that the ALS system would still be controlling the differential degree of freedom.  The values and sign for the 1/sqrt(trans) signals came from a transfer function of dividing the spectra of each error signal and noting the relative gain and sign.

After we swapped the error signals, we realized that we had to remove the offset from the XARM servo, which is why we should have put the offsets elsewhere in the first place.

Then, Koji took a spectrum, which is attached to this entry.  We note that the ALS signals are strongly correlated, and mostly common. 


To Do List

Going forward, we need to figure out what is going on with the PRMI, and why we're having trouble keeping lock.

We need to redo the PRCL ASC servo, with the anti-oplev trick that Rana mentioned a week or two ago.

We need to investigate the degeneracy of REFL165, now that Q's simulation doesn't justify / explain it. 

Attachment 1: common_diff_ALS.pdf
common_diff_ALS.pdf
  9658   Wed Feb 19 18:21:33 2014 manasaUpdateLSCScripts for ALS modified

Quote:

We need to change several scripts for use with the new ALS-in-the-LSC paradigm:

* Watch arms (to turn off ALS if we lose the beatnote, before pushing optics too hard)

* Find IR resonance

* Offset from resonance

None of these should be difficult, just changing the filter bank names to match the new ones (ex. LSC-XARM rather than ALS-XARM, and LSC-ALSX rather than ALS-OFFSETTER1). 

So far, I have changed the "find resonance" script (ALSfindIRresonance.py).  I believe, in principle, to first order, that my modifications should work, however I have not yet tested the script.  So.  If you use it, watch the output of the script and ensure it's doing what it ought.  I'll check it after the lunch meeting and update this log entry.  (I changed the name of the "OFSFILT" variable, line 26, and also modified line 114.  Both of those lines have comments on how to revert the changes).

I have also changed the "offset from resonance" script (ALSchangeOffset.py).  Again, since I'm not locking right now, I have not tested this script either.  So, pay attention if you need to use it, before I check it.  (I changed the name of the OFSFILT variable, and the check which arm logic around line 37.  Again, both of those lines have comments on how to revert the changes.)

Watch arms script (ALSdown.py) has been modified and now watches the LSC-$ARM filter module instead of the ALS-$ARM filter module. Threshold has been kept the same +/-5000 counts to the ETM suspensions. The script has been tested and works just fine. It exists in the same place scripts/ALS/.

Jenne's modified versions of ALSfindResonance.py and ALSchangeOffset.py were tested and work just fine.

  9657   Wed Feb 19 16:42:08 2014 ericqUpdateLSCSome Simulation Efforts

Disregard previous ELOGs, I had the PRC locked on carrier 

Locked on the sideband, the MICH / PRCL angle is much less sensitive to the PRC length, and shouldn't in fact be as degenerate as we've seen in reality. 

SBLOCK_PRMISensingAsIs.pdfSBLOCK_MICHvPRCLangle.pdf

So, my simulations no longer provide any reason for the 3F signals to be so degenerate. 

  9656   Wed Feb 19 14:14:46 2014 ericqUpdateLSCSome Simulation Efforts

 Q EDIT: THIS IS WRONG. I LOCKED PRC ON THE CARRIER

Koji noted oddities in the sensing matrix results I had gotten; namely that the plots showed REFL33 not changing at all, when we know for a fact that this should not be the case. 

Gabriele lent his eyes to my code, and came up with the idea that the modulation depths I was using were maybe not ideal (.1 for both 11 and 55). This affects REFL33 in that it is not simply Carrier * 33Mhz + 11Mhz * -22Mhz but also 22MHz * 55MHz, etc. 

I got more realistic values from Jenne (0.19 for 11MHz and .26 for 55Mhz) and re-ran the code, with more realistic results. The behavior for 165 has remained the same, but the other signals are more well behaved. 

Moral of the story: the modulation depths affect the 3f signals in a complicated way.

PRMISensingAsIs.pdf

PRMISensingCoinc.pdf

MICHvPRCLangle.pdf

 

 

  9655   Wed Feb 19 11:45:12 2014 JenneUpdateLSCScripts for ALS being modified

We need to change several scripts for use with the new ALS-in-the-LSC paradigm:

* Watch arms (to turn off ALS if we lose the beatnote, before pushing optics too hard)

* Find IR resonance

* Offset from resonance

None of these should be difficult, just changing the filter bank names to match the new ones (ex. LSC-XARM rather than ALS-XARM, and LSC-ALSX rather than ALS-OFFSETTER1). 

So far, I have changed the "find resonance" script (ALSfindIRresonance.py).  I believe, in principle, to first order, that my modifications should work, however I have not yet tested the script.  So.  If you use it, watch the output of the script and ensure it's doing what it ought.  I'll check it after the lunch meeting and update this log entry.  (I changed the name of the "OFSFILT" variable, line 26, and also modified line 114.  Both of those lines have comments on how to revert the changes).

I have also changed the "offset from resonance" script (ALSchangeOffset.py).  Again, since I'm not locking right now, I have not tested this script either.  So, pay attention if you need to use it, before I check it.  (I changed the name of the OFSFILT variable, and the check which arm logic around line 37.  Again, both of those lines have comments on how to revert the changes.)

  9654   Wed Feb 19 11:00:16 2014 ericqUpdateLSCSome Simulation Efforts

 Q EDIT: THIS IS WRONG. I LOCKED PRC ON THE CARRIER

 As Koji measured the other day: MICH and PRCL seem very degenerate in the 3f REFL PDs. 

I'm using this as a motivation to do some simulation in MIST and try to understand the best way to implement the 3F locking scheme. Hopefully my thinking below isn't nonsense...

First, I modeled the PRC with no arm cavities and the estimated cavity length I got with the PRM kick measurement, and looked at the REFL sensing matrix.

PRMISensingAsIs.pdf

This agrees with the observed degeneracy. I then modeled the case of the PRC length that gives coincident SB resonance, again with no arm cavities.

PRMISensingCoinc.pdf

Now there is good separation in REFL165. (REFL33 still looks pretty degenerate, however). This raised the question, "What does the angle between MICH and PRCL in REFL165 do as a function of macroscopic PRC length?" 

MICHvPRCLangle.pdf

  • We see ~90 degrees at coincident resonance
  • Shortening the cavity, which we did to account for the arms, quickly shrinks the angle
  • Presuming we moved to make the cavity 4cm shorter implies we had ~45 degrees between MICH and PRCL in REFL165 before the move. (Is this consistent with earlier observations?)

To me, this implies that locking the PRC on 3F from scratch won't be simple. However, the whole point of the PRC length choice is to have coincident SB resonance when the arms are resonating.

So: even if we're not spot on, we should be relatively close to the PRC length where having arms resonant gives us simultaneously resonant upper and lower sidebands, where MICH and PRCL should be orthogonal-ish. I.e. building up a little bit of IR power in the arms may start to break the degeneracy, perhaps allowing us to switch from 1F to 3F locking, and then continue reducing the CARM offset. 

So, I ultimately want to model the effect of arm power buildup on the angle between MICH and PRCL in the 3f PDs. This is what I'm currently working on. 

So far, I have reproduced some of the RC modeling results on the wiki to make sure I model the arms correctly. (I get 37.7949 m as the ideal arm length for a modulation freq of 11.066134 MHz vs. 37.7974m for 11.065399 MHz as stated on the wiki). Next, I will confirm the desired PRC length that accounts for the arms, and then look at the MICH vs PRCL angle in the REFL PDs as a function of arm power or detuning. 

ArmLengthChoice.pdf

  9653   Wed Feb 19 08:07:01 2014 steveUpdatesafetyannual laser safety glasses check

 

 All 40m laser safety glasses are cleaned and measured this morning.  Bring your own safety glasses if you have to enter the 40m IFO room.

 

Glasses were washed in 1% Liquinox water solution and  their transmission measured at  165 mW,  2 mm OD beam of 1064 nm

Attachment 1: AA2014checked.jpg
AA2014checked.jpg
Attachment 2: 2014check.jpg
2014check.jpg
  9652   Wed Feb 19 03:07:22 2014 JenneUpdateLSCALS locked with LSC!

No more progress tonight.  I am still unable to lock the ALS using the regular LSC filters.  I went back to putting the ALS filters into the LSC filter banks, and locked both arms with ALS, and found their IR resonances. I then held them off resonance, and tried to lock PRMI with REFL 55 I&Q, with no success.  Just before locking the arms, I had redone the whole IFO alignment (lock arms in IR, ASS, lock and align MICH, lock and align PRMI), and the PRMI was flashing very nicely.  I'm not sure why I wasn't able to catch lock, except that perhaps 3 or 6 ALS offset counts isn't far enough away from the IR resonance to make the 1f signals happy. The MC lost lock, which I then took as a sign that it's time to go home. (I was hoping to do a quick PRMI + 2arms, and see that we don't lose PRMI lock.  I was going to catch lock with REFL55, then transition to REFL33, although if I had thought about it before the MC lost lock, I would have tried just catching lock with REFL33).

I restored the regular LSC filters for the X and Y arms, and locked the arms in IR just to make sure it's all honkey-dory.  Which, it's not quite.  I don't know why, but right now, neither arm wants its boost (FM9) enabled.  It's part of the restore script that FM9 is triggered along with the rest of the filters, but even if I turn on the filters manually, I can turn on all but FM9, and then when I turn on the boost, the arm falls out of lock. Same behavior for both arms.  Anyhow, they lock, and they seem okay modulo the boost not being able to engage.

  9651   Wed Feb 19 01:33:03 2014 JenneUpdateLSCALS locked with LSC!

I am also not able to lock the ALS using the 'regular' LSC filters.  To figure out what filters were doing what, I made several comparison plots from Foton.

The first one is the progression of ALS locking, using the filters from ALS-XARM.  FM5 is always engaged, then FMs 2, 3, 6, 7, and 8, and finally FM 10 (the low frequency boost) is engaged.

ALS_XARM_LockingFilters.pdf

The next plot is a comparison between the ALS version of the filters, and the LSC-XARM equivalents. 

ALSvsLSC_AllLockingFilters.pdf

Finally, just so I remember which LSC filters do what, I made an equivalent of the first plot, but for the LSC filters.

LSC_XARM_LockingFilters.pdf

When I try to lock the Xarm ALS using the regular LSC filters, I'm getting an oscillation somewhere, that grows and eventually knocks me out of lock.  It looks from dataviewer to be in the ~few Hz range, but it's hard to see it in DTT, since I don't stay locked all that long once the oscillation starts.  (If I catch it, I can back off the gain and turn off the servo without losing lock, but if I don't turn off the servo, I inevitably push the ETM too hard and lose green lock to the arm.)  I tried engaging the 3.2 Hz resonant gain filter, and it just makes things oscillate sooner, so that's not a solution with the current filter designs. 

Also, I'm not able to lock the IR using the ALS version of the XARM filters.  I'll have to meditate more on the situation, but the filters seem to be different enough that there's no crossover at this point.

  9650   Wed Feb 19 00:35:23 2014 KojiUpdateLSCALS locked with LSC!

Great. I indeed disabled all of the triggers and the normalization during my trial but in vain.
So I'm curious this is actually because of the filter shape or not.

  9649   Tue Feb 18 23:55:33 2014 JenneUpdateLSCALS locked with LSC!

I'm really excited, so I'm posting this, even though I'm still working:

I currently have ALS locked using the LSC system, and have (by hand, coarsely) found IR resonance!  Hooray!

I looked at my error signals, as well as LSC-XARM_IN1 with dataviewer, and noticed that the XARM_IN1 signal was crazy when I was using the ALS signal as the error.  I soon realized that this is because there was a non-zero element in the power normalization matrix, and I'm overriding the trigger.  So, I was trying to divide by zero, and was getting crazy numbers.  After zeroing the power normalization matrix element for the Xarm, the XARM_IN1 signal matched the ALSX_OUT, and I was easily able to acquire lock.

I had already re-transferred over the ALS versions of the filters, so that's what I'm using right now.  Next up (on a 5 minute time-scale) is trying to acquire lock using the regular LSC filters. 

Oh, also, something I hadn't thought of before dinner:  I am setting the offset of the ALSX filter bank such that the output is centered around zero, so that I can lock, since these are not AC coupled servos.

  9648   Tue Feb 18 23:27:14 2014 JenneUpdateLSCALS not locking with LSC

It looks like its somehow a discrepancy between the TFs of each error signal, because features are similar, and present, in both error signals.

ALSX_POXvsBeatnote_withEXCtfs.pdf

  9647   Tue Feb 18 20:31:29 2014 KojiUpdateLSCALS not locking with LSC

Hmm. Wierd. Can you look at the TFs between ETMX-EXC and the error signals so that we can identify which one has these structures.

  9646   Tue Feb 18 18:52:08 2014 JenneUpdateLSCALS not locking with LSC

Koji mentioned to me (and elogged) that he was unsuccessful locking the ALS using the LSC servos.  He suggested I look into this.

So, rather than just looking at the transfer function between POX or POY and the green beatnotes at a single frequency, I did a whole transfer function.  The point was to see if the TF is flat, and if we get any significant phase lag in the transfer from c1als to c1lsc.  (c1als is running on the IOO machine, so an RFM connection is involved in getting it over to the LSC machine.)

In the first figure, I have plotted POX vs. Beatnote_PHASE_OUT (ALS error signal, still in the c1als model), and POX vs. ALSX_IN1 (the ALS error signal, after transfer over to the c1lsc model).  You can see that we have a little phase lead in the blue transfer function, and fairly significant phase lag in the red (red is after transfer over to the lsc model).  In the grand scheme of things, the magnitude is fairly flat, however that is not perfectly true - the peaks seen near 50 Hz and 300Hz are repeatable.  The relative phase lag between the "BEATX" version of the signal in the ALS model, and the "ALSX" version of the signal in the LSC model is 15 degrees at 200 Hz, which corresponds to 33 usec.   

ALSX_POXvsBeatnote.pdf

The second figure is the same as the first, except for the Yarm.  The relative phase lag between the ALS version of the error signal and the LSC version is 16 degrees at 200 Hz, which is about 35 usec.

ALSY_POYvsBeatnote.pdf

As a side note, before trying any ALS locking, I took a spectrum of the beatnote (in the ALS model) while the arms were locked with IR:

BeatNoteSpectra_ArmsLockedWithIR_28Feb2014.pdf

To check things, I made sure that I could lock the Xarm ALS using the old ALS system - I was able to do so.  (Has someone put the "watch" script as a constantly-on thing?  It's kind of nice not to have to turn it on, although we'll need to change it to turn off the LSC versions of the servos eventually). 

Then, I tried locking the Xarm using the LSC system (using only FM5 of the regular LSC-XARM filter bank).  Like Koji, I was not able to acquire lock.  As a next step, I copied all of the LSC-XARM filters into an empty filter module, LSC-XXXDC (the first one on the list underneath LSC-XARM), and copied over the ALS Xarm filters to the LSC Xarm filter bank.  I then tried to acquire lock, but am unable to get it to stay.  Using the ALS system, when you put in a small gain, the beatnote starts to settle down, and as you increase the gain, the beatnote stops moving (as seen on the spectrum analyzer) almost completely.  However, using the LSC system, the beatnote never really stops moving or settles down.  And if I increase the gain, I push the ETM hard enough that I lose green lock.  I have put the regular LSC filters back for now.

Here is a plot from Foton comparing the FM5 filter modules from the LSC-XARM (regular IR locking) and the ALS-XARM servo.  They are pretty different, and have 10 degrees of phase difference at 200 Hz, because 2 of the 3 poles are complex in the LSC version, while the ALS version is just a single real pole.

ALSvsLSC_LockingFilters.pdf

Anyhow, I am declaring it to be dinnertime, and I plan to return in a few hours. Since I put the regular LSC filters back (since I'm going to have to realign after dinner anyway), the IFO should be in its nominal state if anyone wants to come in and play with it.

  9645   Tue Feb 18 14:28:15 2014 JenneUpdateIOOMC unstable - centering spots helped

As we've been seeing a bit lately, the MC will be locked happily for several hours, but then it will start misbehaving. 

Today, I measured the spots on the MC mirrors, and found that the MC2 spot was quite far off in yaw (about -3.5 cm).  I recentered the MC2 spot, and then (with the MCWFS on), moved MC1 and 3 until their WFS outputs were close to zero (they had gone up to 100+).  In the ~15 minutes since doing that, the MC refl signal is not oscillating like it was, the transmission is up, and the MC has not unlocked. 

To reiterate, I did not touch any settings of anything, except the alignment of the MC mirrors to center the MC2 spot, and then offload the WFS.  Next time the MC starts acting up, we should measure the spots, and roughly center them, before messing with any other settings.  Note however, that this is a ~10 minute procedure (including the fact that one spot measurement takes a little less than 5 minutes).  This need not be a several hour endeavour. 

  9644   Tue Feb 18 10:38:55 2014 SteveUpdateVACRGA scan at day 13

 All normal.

 

Attachment 1: pd77m13d.png
pd77m13d.png
Attachment 2: RGAscan13d.png
RGAscan13d.png
  9643   Tue Feb 18 10:10:07 2014 SteveUpdatePEM floor cleaning under racks

Quote:

 Keven, Steve

 

The floor was cleaned under the east arm tube with hand held wet towel. We moved staff around and mopped. I did at the bottom of rack  1Y1, 2  and 3. 

Last week we did the south arm tube floor.

Next week we 'd like to clean under rack 1X1,2,3, 4, 5, 6 and 7 

 Keven, Steve

1Y4, 1X1,2,3,4 & 5 instrument racks floor space were cleaned.

  9642   Mon Feb 17 20:35:19 2014 ericqUpdateElectronicsTransmon QPD whitening

My apologies for all of that crap I left at the Y-end... I cleaned the rest of it up today. 

I took transfer functions of the four ETMY QPD whitening channels today. (Attempted the ETMX ones too, but had troubles driving the board; detailed below). I've attached a zip with the DTT xml files for the cases of no whitening / 1 whitening stage / both whitening stages engaged. Here's a plot of both whitening stages engaged. 

qpdY2whit.pdf

 

Given the way I measured, the DAC output anti-imaging is in the TFs as well. ( This is a D000186 board; with something like a 4th order elliptic LP, but I need to look at the board / fit the TF to see the parameters, there are different revisions with different filter shapes.) 

The c1scy model had excitation blocks on some of the unused DAC channels (C1:SCY-XXX_CHAN9 etc.), but these were in the second DAC output connection, and not cabled up. However, the 8th channel on the DAC had no connection in the simulink model, so I added another excitation block there (C1:SCY-XXX_CHAN8), and used the anti-imaging front panel lemo connector to drive the input of the whitening board. 

I also added a similar channel to the SCX model, but no data would show up in the channel as viewed by data viewer (though the channel name was black), or in analog world. There's the additional weirdness that the SCY excitation channels show up under SCX in DTT and awggui... I'm not entirely sure what's going on here.

I still need to look at the noise, and peek inside the boards, to check for homemade modifications and see if there are bad things like thick film resistors that may be spoiling the noise performance...

Attachment 2: ETMY_QPD_whitening.zip
  9641   Sun Feb 16 17:40:11 2014 KojiUpdateLSCFriday Night ALS

I wanted to try common/differential ALS Friday evening. I tried ALS using the LSC servo but this was not successfull.
The usual ALS servo in the ALS model works without problem. So this might be coming from the shape of the servo filter.
The ALS one has 1:1000 filter but the LSC one has 10:3000. Or is there any problem in the signal transfer between
ALS and LSC???


- MC:
Slow offset -0.302V

- IR:
TRX=1.18 / TRY=1.14, XARM Servo gain = 0.25 / YARM Servo gain = 0.10

- Green Xarm:
GTRX without PSL green 0.562 / with PSL green 0.652 -> improved upto 0.78 by ASX and tweaking of PZTs
Beat note found at SLOW OFFSET +15525
Set the beat note as +SLOW OFFSET gives +BEAT FREQ

- Green Yarm:
GTRY without PSL green 0.717 / with PSL green 1.340
Beat note found at SLOW OFFSET -10415
Set the beat note as +SLOW OFFSET gives -BEAT FREQ

- BEAT X -10dBm on the RF analyzer@42.5MHz / Phase tracker Qout = 2300 => Phase tracking loop gain 80 (Theoretical UGF = 2300/180*Pi*80 = 3.2kHz)
- BEAT Y -22dBm on the RF analyzer@69.0MHz / Phase tracker Qout = 400 => Phase tracking loop gain 300 (Theoretical UGF = 2.1kHz)

Transfer function between ALSX/Y and POX/Y11I @560Hz excitation of ETMX
POX11I/ALSX = 54.7dB (~0deg)
POY11I/ALSY = 64.5dB (~180deg)

POX11I calibration:

ALSX[cnt]*19230[Hz/cnt] = POX11I[cnt]/10^(54.7/20)*19230[Hz/cnt]
= 35.4 [Hz/cnt] POX11I [cnt] (Hz in green frequency)

35.4 [Hz/cnt]/(2.99792458e8/532e-9 [Hz]) * 37.8 [m] = 2.37e-12 [m/cnt] => 4.2e11 [cnt/m] (c.f. Ayaka's number in ELOG #7738 6.7e11 cnt/m)

POY11I calibration:

ALSY[cnt]*19230[Hz/cnt] = POY11I[cnt]/10^(64.5/20)*19230[Hz/cnt]
= 11.5 [Hz/cnt] POX11I [cnt] (Hz in green frequency)

11.5 [Hz/cnt]/(2.99792458e8/532e-9 [Hz]) * 37.8 [m] = 7.71e-13 [m/cnt] => 1.3e12 [cnt/m] (c.f. Ayaka's number in ELOG #7738 9.5e11 cnt/m)

  9640   Fri Feb 14 21:03:13 2014 KojiUpdateGeneralY end "BS"

As I didn't have the green laser PZT feedback for the laser temp control, I went to the yend to check out what's the situation.

I found horrible and disgusting "remnants".

WHAT ARE THESE BSs AT THE Y END?

- The table enclosure was left open

- A (hacky) DB25 cable with clips was blocking the corridor and I was about to trip with the cable.

- This DB25 cable went to the table without going through the air tight feedthrough that is designed for this purpose.

- An SR560 (presumably for the openloop TF measurement) was left inserted in the loop with entangled cables connected to the servo box.

- Of course the laser PZT out mon was left unplugged.

Even after cleaning these cables (a bit), the end setups (including the X end too) are too amature.
Everything is so hacky. We should not allow ourselves to construct this level of setup everytime
we work on any system. This just adds more and more mysteries and eventually we can't handle
the complexity.

  9639   Fri Feb 14 12:39:02 2014 JenneUpdatesafetySmoke detector cleaned

Facilities just came by and cleaned the smoke detector that is above Steve's desk.  It's next to an air vent, so I guess it collects dust more than a "typical" smoke detector.

  9638   Fri Feb 14 02:33:09 2014 manasaUpdateGeneralMy IFO time summary

MC tuning

Although the morning MC tuning looked stable, Koji pointed out that the MC_REFL_OFFSET was changed from its nominal value.

The offset was reset and this caused drift in the MC_TRANS_SUM.

To fix this:

- disabled the WFS servo

- aligned MC using MC1 and MC3

- centered beam on the MC_REFL

- reset WFS offsets

- locked MC

MC looks happy now.

__________________________________________________________________________________________________________________

ALS locking

ALS is in a very different state from a couple of days ago when we could successfully lock the arms and scan.

The green alignment to the arms had drifted.

PSL green alignment on the PSL table was off. The PSL green was not even on the steering mirror. Did anyone work around the PSL table in the last couple of days?

After aligning and finding the beat note, I found the ALS servo very noisy. The error signal had 10 times more rms noise than what was achieved earlier this week and there were some new 60Hz peaks as well.

Overall, we could not do any PRMI+ALS arms today

 

 

 

 

  9637   Fri Feb 14 02:09:55 2014 ericqUpdateElectronicsTransmon QPD whitening

 [Quick post, will follow up with further detail later. Excuse my sleepy ELOG writing]

Goal: Check out the transmon QPD signal chain; see if whitening works. Assess noise for 1/sqrt(TRX/Y) use. 

First impression: Whitening would not switch on when toggling the de-whitening. The front monitors on the whitening boards are misleading; they are taken a few stages before the real output. ADC noise was by far the limiting noise source. 

I updated the binary logic in the c1scx and c1scy to actually make the binary IO module output some bits. 

After consulting a secret wiring diagram on the wiki, not linked on the rack information page (here), I worked out which bits correspond to the bypass switches in the whitening board ( a fairly modified D990399, with some notes here)

Now, FM1 and FM2 (dewhitening filters on the ETM QPD quadrants) trigger the corresponding whitening in the boards. Here's a quick TF I took of the quadrant 1 board at ETMY. (I should take a whitening+dewhitening TF too, and post it here...)

qpdWhitening.pdf

Seems to roughly work. Some features may be due to non-accounted for elements in the anti-imaging of the DAC channels I used for the excitation, or such things. The board likely needs some attention, and at least a survey of what is there. 

I also need to take dark noise data, and convert into the equivalent displacement noise in the 1/sqrt(TRX/Y) error signals. For the no-whitening ADC noise, I estimated ~1pm RMS noise on a 38pm linewidth of PRFPMI arms. 

  9636   Fri Feb 14 00:58:41 2014 JenneUpdateLSCThoughts on Transition to IR

[Koji, Jenne, EricQ, Manasa]

We had a short discussion this evening about what our game plan should be for transitioning from using the ALS system to IR-generated error signals. 


The most fundamental piece is that we want to, instead of having a completely separate ALS locking system, integrate the ALS into the LSC.  Some time ago, Koji did most of the structural changes to the LSC model (elog 9430), and exposed those changes on the LSC screen (elog 9449).  Tonight, I have thrown together a new ALS screen, which should eventually replace our current ALS screen.  My goal is to retain all the functionality of the old screen, but instead use the LSC-version of the error signals, so that it's smoother for our transition to IR.  Here is a screenshot of my new screen:

Screenshot-Untitled_Window-1.png

You will notice that there are several white blocks in the center of the screen. From our discussion this evening, it sounds like we may want to add 4 more locking servo paths to the LSC (ALS for each individual arm, and then ALS for CARM and DARM signals).  The reason these should be separate is that the ALS and the "regular" PDH signals have different noise characteristics, so we will want different servo shapes.  I am proposing to add these 4 new servo blocks to the c1lsc model.  If I don't hear an objection, I'll do this on Monday during the day, unless someone else beats me to it.  The names for these filter modules should be C1:LSC-ALS_XARM, C1:LSC-ALS_YARM, C1:LSC-ALS_DARM and C1:LSC_ALS_CARM.  This will add new rows to the input matrix, and new columns to the output matrix, so the LSC screen will need to be modified to reflect all of these changes.  The new ALS screen should automatically work, although the icons for the input and output matrices will need to be updated. 

The other major difference between this new paradigm and the old, is the place of the offset in the path.  Formerly, we had auxiliary filter banks, and the summation was done by entering multiple values in the ALS input matrix.  Now, since there is a filter bank in the c1lsc model for each of the ALS signals precisely where we want to add our offsets, and I don't expect us to need to put any filters into those filter modules, I have used the offset and TRAMP of those filter banks for the offsets.  Also, you can access the offset value, and the ramp time, as well as the "clear history" button for the phase tracker, all from the main screen, which should help reduce the number of different screens we need to have open at once when locking with ALS.  Anyhow, the actual point where the offset is added has not changed, just the way it happens has. 

When we make the move to using the ALS in the LSC, we'll also need to make sure our "watch arm" and "scan arm" scripts are updated appropriately.

As an intermediary locking step, we want to try to use the ALS system to actuate in a CARM and DARM way, not XARM and YARM.  We will transition from using each ALS signal to feed back to its own ETM, to having DARM feed back to the ETMs, and CARM feed back to MC2.  We may want to break this into smaller steps, first lock the arms to the beatnotes, then find the IR resonance points.  Transition to CARM and DARM feedback, but only using the ETMs.  After we've done that, then we can switch to actuating on MC2.  If we do this, then we'll be using the MC to reduce the CARM offset.

Once we can do this, and are able to reduce the CARM offset, we want to switch CARM over to a combination of the 1/sqrt(transmission) signals.  The CARM loop has a tighter noise requirement, so we can do this, but leave DARM locked to the beatnotes for a while.

After continuing to reduce the CARM offset, we will switch CARM over to one of the RF PDs, for its final low-noise state. 

We'll then do a quick swap of the DARM error signal to the AS port (maybe around the same time as CARM goes over to a PDH signal, before the CARM offset is zero?). 

During all of this, we hope that the vertex has stayed locked. If our 3f sensing matrix elements are totally degenerate when the arms are out of resonance, then we may need to acquire lock using REFL 1f signals, and as we approach the delicate point in the CARM offset reduction, move to 3f signals, and then move back to 1f signals after the arm reflection has done its phase flip.  Either way, we'll have to move from 3f to 1f for the final state.

  9635   Thu Feb 13 22:27:54 2014 JenneUpdateLSCPRMI locked on REFL 33 I&Q

[Jenne, Koji]

I was able to get the PRMI locked on REFL33 I&Q, but it wasn't overly stable, since there is so little separation between the MICH and PRCL signals in that PD. 

We have already adjusted the phase to maximize PRCL in the I-phase.  Since MICH is ~45 degrees separated from PRCL, there is some projection of MICH in the I-phase, and some in the Q-phase. 

To remove this MICH component, I locked the PRMI on REFL55, and drove MICH.  I looked at REFL33I at the CARM filter bank input (as just a dummy location to get a signal into DTT).  I then added REFL33Q to the CARM row of the input matrix, to try to get the MICH line minimized.  I then used these values for PRCL, and used just REFL33Q for MICH, and re-locked the PRMI.  The PRMI was much more stable and happy. 

The input matrix values that I used were:

MICH:  REFL33Q = -0.487, Servo Gain = -20.0

PRCL:  REFL33I = 1.556, REFL33Q = 1.8, Servo Gain = -0.020

 

Some locking notes:

The PRMI is very sensitive to alignment, and the PRM tends to drift away from optimal alignment on a ~1 hour timescale.  When the PRM was not well aligned, it looked like MICH had a locking offset (manifested as non-equally sized blobs at AS).  The MICH offset seemed to go away when we realigned the PRM. 

ELOG V3.1.3-